User name Elena Kh.
Log entry time 01:57:38 on April17,2007
Entry number 199316
keyword=Changed pedestal suppression threshold before run 2559
During the day the pedestal suppression thresholds were changed. Makis Petratos asked me to restore them (->10) for left/right s1, s2 and the shower detector again like they were 24 hours ago.
The plots are from run 2557.